OpenStudy (anonymous):

What is the slope of the line between (-4,4) and (-1-2)? please help i have little time I know this is easy, but I did it and the answer is not one of the choices :( I got -3, but the choices are: 1 ,2, -2, or -1.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (amistre64):

show your work ...

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

if youre trying the formula, then you are most likely like me and have placed something in the wrong spot

OpenStudy (anonymous):

(-4,4) will be (x1,y1) (-1,-2) will be (x2,y2)

OpenStudy (anonymous):

then it will be: -2-4 over -1- (-4) = -6 over 3 = -3

OpenStudy (amistre64):

i do a vertical method, subtracting down the columns (-4,4) -(-1-2) ------- (-3,6) , now slope is defined as y/x 6/-3

OpenStudy (anonymous):

wait what.. i did it so wrong

OpenStudy (anonymous):

you just subtract?

OpenStudy (anonymous):

that's not how we were shown to do it so i dont think i could do it that way..

OpenStudy (amistre64):

my method is to stack one point on top of the other, and subtract down the columns

OpenStudy (anonymous):

that's not the right way to do it though..

OpenStudy (amistre64):

it is the right way to do it ...

OpenStudy (anonymous):

but i usually get it right.. how come this time something went wrong

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

look how i did it ^^

OpenStudy (anonymous):

\[\frac{ -2-4 }{ -1-(-4) } - \frac{ -6 }{3 } = -2\] ohhhhh

OpenStudy (anonymous):

i just put -6 divided by 3 = -3, instead of -2!!!!!!!!!

OpenStudy (anonymous):

ok that's the stupidest mistake ever

OpenStudy (amistre64):

take 2 points: (xo,yo) and (x1, y1) stack them and subtract (its just a method to organize the information) (xo,yo) -(x1, y1) --------- xo-x1, yo-y1 , you still get the same information; now slope is defined as y/x yo - y1 ------- xo - x1

OpenStudy (anonymous):

i just waste 5 mins

OpenStudy (amistre64):

since i tend to place the wrong value into the wrong places .... i organize the data in a more convenient manner that is harder to mess it up